A stamp collection consists of 8-cent, 13-cent, and 18-cent stamps. The number of 8-cent stamps is twice the number of 18-cent stamps. There are 3 more 13-cent stamp than 18-cent stamps. The total value of the stamps in the collection is $3.68. Find the number of 8-cent stamps in this collection.
I tried letting
X = 18-cent stamps
2x = 13-cent stamps
X+3 = 8-cent stamps
There's a similar problem on the site, but I can't make it work for my own problem. I keep getting decimal answers, when I think it should be whole numbers (number of stamps; you can't really have a fraction of a stamp).

Answer by checkley77(12569)

X = 18-cent stamps
2x = 8-cent stamps
X+3 = 13-cent stamps
You had the .08 cent & .13 cent stamps mixed up.
.18x+.08*2x+.13(x+3)=3.68
.18x+.16x+.13x+.39=3.68
.47x=3.68-.39
.47x=3.29
x=3.29/47
x=7 18 cent stamps.
2*7=14 8 cent stamps
7+3=10 13 cent stamps.
Proof:
.18*7+.08*14+.13*10=3.68
1.26+1.12+1.30=3.68
3.68=3.68